OpenStudy (loser66):

Find L1L2 and L2L1 for the given differential operators and determine whether L1L2= L2L1 L1= D +1 L2= D - 2x^2 Please, help

OpenStudy (loser66):

@primeralph

OpenStudy (primeralph):

Just check if they commute.

OpenStudy (loser66):

how?

OpenStudy (loser66):

I got it wrong, since L1 = D +1 so L1 = y' +y and L2= D -2x^2 = y' - 2x^2 y L1L2 = (y'+y)((y' +2x^2 y) is it right?

OpenStudy (loser66):

and then L2L1 are the same, but the answer in the back of the book said that they are not equal :(

OpenStudy (primeralph):

L1L2 means you apply the operation on L2; so that's L2' +L2. So for L1L2y = L1(y'-2x^2 y) = (y'-2x^2 y)' + (y'-2x^2 y)

OpenStudy (primeralph):

Commuting means L1L2 - L2L1 = 0.

OpenStudy (loser66):

so?

OpenStudy (primeralph):

So read what I wrote.

OpenStudy (loser66):

L2'+L2 for L1L2, right? L1' -2x^2L1 for L2L1, right?

OpenStudy (primeralph):

Then you have to apply them on some function and see if they are the same.

OpenStudy (loser66):

got you, thanks a lot

OpenStudy (primeralph):

You're welcome.